40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 25 10:31:33 2012, Sasha, Summary, Simulations, SURF - Week 5 - Summary Screen_Shot_2012-07-24_at_10.46.05_AM.pngScreen_Shot_2012-07-24_at_10.45.43_AM.pngScreen_Shot_2012-07-25_at_10.16.58_AM.pngScreen_Shot_2012-07-25_at_10.25.27_AM.png
    Reply  Wed Jul 25 14:35:45 2012, Sasha, Summary, Simulations, SURF - Week 5 - Summary Screen_Shot_2012-07-25_at_2.20.01_PM.pngScreen_Shot_2012-07-25_at_2.20.15_PM.pngScreen_Shot_2012-07-25_at_2.20.29_PM.pngScreen_Shot_2012-07-25_at_2.23.05_PM.png
Message ID: 7028     Entry time: Wed Jul 25 14:35:45 2012     In reply to: 7022
Author: Sasha 
Type: Summary 
Category: Simulations 
Subject: SURF - Week 5 - Summary 


This week I've been working on refining my simulation and getting it ready to be plugged into the control system. In particular, I've added a first attempt at a PDH control system, matrix conversion from OSEMs to DOF and back, and all necessary DAC/ADC/AA/AI/whitening/dewhitening filters. Most of these work well, but the whitening filters have been giving me trouble. At one point, they were amplifying the signal instead of flatting it out, such that my simulation started outputting NaN (again).

This was wholeheartedly depressing, but switching out the whitening filters for flat ones seemed to make the problem go away, but brought another problem to light. The output to input ratio is minuscule (as in 10^-300/10^243, see Attachment 3 for the resulting bode plot between a force on the suspension pt in the x-direction and my two outputs - error signal and length signal, which is pretty much what you would expect it to be). I suspect that its related to the whitening filter problem (perhaps the dewhitening filter is flattening the signal instead of amplifying?). If that is the case, then switching the whitening/dewhitening filters ought to work. I'll try today and see what happens. The white/dewhite filters together result in a total gain of 1, which is a good fundamental test, but could mean absolutely nothing (i.e. they could both be wrong!). Judging from the fact that we want to flatten out low frequency signal when it goes through the whitening filter, the filters don't look switched (see Attachment 4 for a bode plot of white and dewhite).

The only other source of problems (given that the suspensions/local damping have been debugged extensively throughout this process - though they could bug out in conjunction with the cavity controls?) is the PDH system. However, separating each of the components showed that the error signal generated is not absurd (I haven't tested whether it makes sense or not, but at any rate it doesn't result in an output on the order of 10^-300).

In summary, I've made progress this week, but there is still far to go. Attachment 1 is my simulation from last week, Attachment 2 is my simulation from this week. A talk with Jamie about the "big picture" behind my project helped tremendously.

 Here's a screenshot of what's going on inside the cavity (Attachment 1). The PDH/mixer system outputs 0 for pretty much everything except really high numbers, which is the problem I'm trying to solve now. I assumed that the sidebands were anti-resonant, calculated reflection coefficient F(dL) = Z * 4pi * i/(lambda), where Z = (-r_1 + r_2*(r_1^2 + t_1^2)/(1 - r_1*r_3)), then calculated P_ref = 2*P_s - 4sqrt(P_c*P_s) * Im(F(dL)) * sin(12.5 MHz * t) (this is pictured in Attachment 2), then mixed it with a sin(12.5MHz * t) and low-passed it to get rid of everything but the DC term (this is pictured in Attachment 3), which is the term that then gets whitened/anti-aliased/passed through the loop.


Attachment 1: Screen_Shot_2012-07-25_at_2.20.01_PM.png  16 kB  | Show | Show all
Attachment 2: Screen_Shot_2012-07-25_at_2.20.15_PM.png  7 kB  | Show | Show all
Attachment 3: Screen_Shot_2012-07-25_at_2.20.29_PM.png  15 kB  | Show | Show all
Attachment 4: Screen_Shot_2012-07-25_at_2.23.05_PM.png  9 kB  | Show | Show all
ELOG V3.1.3-